# Galactic transient sources with the Cherenkov Telescope Array

### v1

@@ 5 Synergies with large astronomical facilities
Simultaneous coordinated observations with telescopes and facilities at different wavelengths are crucial for understanding the processes and mechanisms at work in the sources of our interest (see Cherenkov Telescope Array Consortium et al., 2019 for a detailed review). Coordination, response to real-time alerts and target-of-opportunity (ToOs) are key in time domain astrophysics when observing transient events. In the case of Galactic transients, external alerts can trigger ToO observations by CTA of new Galactic events or renewed activity of known sources. The trigger criteria is dependent on the type of source observed, with different time urgency and duration varying in terms of the evolution of the specific phenomena. We recommened CTA Observaroty to participate in networks (such as the Astrophysical Multi-messenger Observatory Network, AMON999https://www.amon.psu.edu/ or Simultaneous Multiwavelenth Astronomy Research in Transients NETwork, SMARTNet Middleton et al., 2017) to increase the chances of coordination and follow-up in a multi-wavelength and multi-messenger context.
@@
The radio band is of general interest for providing information of the non-thermal processes, and radio facilities are key for locating acceleration sites and shocks. Among other discoveries, the recent association of a magnetar and a FRB has proven that magnetars are the engines of at least some FRBs. FRB-like events triggered in radio will be key to search for (short-timescale) VHE counterparts of, at least, some magnetar engines. For that purpose, CTA will need an external trigger from a radio observatory sensitive to millisecond bursts and capable of issuing prompt alerts. Current generation of telescopes like CHIME, with a large field of view and collecting area, have proven to be excellent detectors for FRB events. Similarly, MeerKat (and future SKA), with its MeerTRAP project to continuously search and detect this type of millisecond-duration events, will provide the needed trigger. Other facilities such ALMA will share important synergies with CTA.
@@
In the optical and infrared domain prompt reaction would be facilitated by coordinating with external observatories within time zones close to those of CTA sites. In the case of CTA North, it should be pursuing synergies mainly with La Palma (and CAHA) telescopes in Spain, while in the CTA South the most appropriate choice would be the different ESO telescopes and other large facilities in Chile, such as the Rubin Observatory Legacy Survey of Space and Time (LSST), which will be key for time-domain astronomy (see Hambleton et al., 2022). The telescope array BlackGEM 101010https://astro.ru.nl/blackgem/ and the he Gravitational-wave Optical Transient Observer (GOTO)111111https://goto-observatory.org/ will perform optical counterpart identification of gravitational wave events, providing wide field transient discovery in the GW era. In both hemispheres the availability of imaging and spectroscopic instrumentation does largely fulfill the observational needs required for coordinated campaigns. Specific target of opportunity proposals could be eventually placed by the CTA community. In parallel, small optical telescopes either on-site or operated by nearby CTA institutions could also play a key role in this effort if the targets are bright enough as in the case of novae or gamma-ray binaries. This could serve to mitigate the often high time pressure on large telescopes. Additionally, support optical telescopes would ensure a fast follow-up on transients and would provide sufficient coverage in the case of bright sources.
@@
Future space missions will allow to study the X-ray domain with improved capabilities. The Space Variable Objects Monitor (SVOM) is expected to be launched in mid 2024 to study GRBs in the 4 - 150 keV range. Althought its science is focused in GRB detection, it is likely that it will also detect outbursts or GRB-like events of Galactic origin, i.e., magnetar giant flares. The Advanced Telescope for High-ENergy Astrophysics (Athena) is scheduled for launch in mid 2030s. It will operate in the energy range of 0.2–12 keV. These and other future observatories will be key in Galactic transient detection. As an example, it was seen that changes in the hardness of the X-ray emission of magnetars could be an indicator of an incoming FRB-like event (Mereghetti et al., 2020), which might allow to trigger observations by CTA if confirmed.
@@
The high-energy band (E100 MeV) is currently explored by the Fermi-LAT satellite. Triggers from Fermi-LAT have been important for the detection of AGNs and novae in the VHE domain, among others. MeV-GeV information provided with simultaneous observations with Fermi is also important to disentangle between hadronic and leptonic processes. If Fermi lifetime extends through (at least partially) CTA lifetime, it will provide the needed coverage and alerts to certain observations/triggers. Just recently, the AGILE satellite, that also operated in this energy range, ceased its operations and re-entered the atmosphere (Tavani et al., 2024). For the moment, there is no obvious successor to Fermi or AGILE in the same energy domain, although some missions could take over. A successor will be crucial to fully exploit the full potential of gamma-ray astronomy. The High Energy cosmic-Radiation Detection facility (HERD) is expected to be operational from 2026 on board of the Chinese Space Station, detecting gamma rays above 0.5 GeV.
@@
Finally, in a multi-messenger context, CTA will be able to search for the electromagnetic counterpart of GWs generated in the Milky Way, such as white dwarf binary mergers. The Laser Interferometer Space Antenna (LISA) mission will be the first GW observatory on space (scheduled launch on 2030) and will key to study white dwarf mergers (Lamberts et al., 2019; Georgousi et al., 2023). CTA could potentially follow-up transient alerts to search for a putative counterpart. Finally, neutrino alerts will also be crucial in the case some events such as e.g. Galactic core-collapse supernovae.
@@ 6 Summary and conclusions
We have studied the capabilities of the forthcoming CTA Observatory to detect transient and variable emission from Galactic sources of different nature. CTA will be able to discover new transients with not significant degradation in the sensitivity, with a maximum of 15% in the crowded inner regions of the Galaxy when overlapping with strong emitters. Similarly, in order to detect variability from dim systems, our simulations have shown that sources with a photon flux ph cm-2 s-1 will require h of observations to detect this variability. For sources with fluxes above this threshold, only h are required. In the case of strong sources ( ph cm-2 s-1), short exposures h are required, implying that low variations in the flux can be detectable from bright sources.
For the case of generic transient sources with the fluxes ph cm-2 s-1 TeV-1 CTA will not detect any source in about one hour observation time, while for those with fluxes ph cm-2 s-1 TeV-1 in an uncrowded region, CTA will be able to detect of sources within just 1 hour of observation for both arrays for the different zenith angles and configurations of the geomagnetic field used in this study.
The unique sensitivity at short timescales together with the fast slewing capabilities of the LSTs and the aforementioned capacities of CTA will allow the detection and discovery of a variety of sources of different nature, according to our simulations:
@@
CTA will detect VHE from microquasars and from the interaction between their jets and the surrounding environment. Our simulations show that CTA will likely detect both transient and persistent emission from the massive microquasars Cyg X-1 and Cyg X-3. CTA will also significantly detect SS 433, both the interaction regions between the jet and the surrounding nebula and possibly the central source. In the case of LMXBs, CTA will detect outbursts within few tens of minutes from a nearby source (4 kpc) with relatively small inclination angle (30∘). Longer exposure times are required for LMXBs with larger angles.
@@
We tested the case of tMSPs, concluding that CTA will need long integration times (50 h) to be able to detect the possible emission of tMSPs when they are in the LMXB state. These systems could be detected during a transition from RSMP to LMXB if an additional VHE component is present, which could provide crucial information on particle interaction.
@@
Flaring emission from the Crab Nebula will be best detected by CTA (or LST sub-array) at low energies (E200 GeV) in less than 1 h. In the TeV regime, integration times of 10 h will be needed, specially for the detection of dimmer flares.
@@
In the case of novae, CTA will be able to detect close-by novae of both classical and symbiotic nature. As an example of the only VHE novae known to date, our simulations reveal that CTA will detect the symbiotic recurrent nova RS Oph with high significance in only 30 min, allowing for a detailed measurement, and therefore detailed modeling, of its SED from energies as low as 20 GeV. Combined with multi-wavelength observations, the temporal and spectral analyses of CTA observation would improve our understanding of the acceleration processes in novae.
Regarding sources of different nature from the aforementioned, we can expect CTA to detect emission from magnetars during a giant flare and even likely during intermediate flares associated with an FRB. Other possible transient events are flares from SFXTs, for which a detection would definitely identify SFXTs as VHE emitters. Other variable VHE candidates are runaway stars and young stellar objects. Serendipitous discoveries are also likely while performing, i.e., surveys.
Simultaneous multi-wavelength observations will be crucial to maximize the scientific output of the CTA observatory. Synergies with the next generation of astronomical facilities will play a key role both for triggering observations with CTA and to perform follow-ups. In a multi-messenger context, we also expect CTA to be able to follow-up GW events of Galactic origin, with prospects for detecting a TeV counterpart.
@@ Acknowledgements
,121212https://www.gammapy.org a community-developed core Python package for TeV gamma-ray astronomy.
@@
The official CTA observatory IRFs (Maier et al., 2023) used in this manuscript are available in: prod3b-v2
@@
A. Aguasca-Cabot, simulations and drafting of the novae subsection; M. Chernyakova, simulations and drafting of the SS 433 and microquasars subsections; D. Kantzas, simulations and drafting of the LMXBs subsection and modeling of Cygnus X-1 microquasar; A. López-Oramas, coordinator and editor of the manuscript, general drafting; P. L. Luque-Escamilla, drafting of optical and infrared synergies and other expected variable gamma-ray sources at TeV subsections; S. Markoff, simulations and drafting of the LMXBs subsection and modeling of Cygnus X-1 microquasar; J. Martí, drafting of optical and infrared synergies and other expected variable gamma-ray sources at TeV subsections; D. de Martino, internal review of the manuscript; S. McKeague, simulations and drafting of the SS 433 and microquasars subsections; S. Mereghetti, drafting of the magnetar subsection; E. Mestre, simulations and drafting of the Crab Nebula flares subsection; A. Mitchell, simulations and drafting of the novae subsection; E. de Oña-Wilhelmi, simulations and drafting of the Crab Nebula flares and the novae subsections; G. Piano, simulations and drafting of the Cygnus X-3, Cygnus X-1, V404 Cyg and microquasars subsections; P. Romano, early perspectives with CTA, SFXT discussion and internal review of the manuscript; A. J. Ruiter, simulations and drafting of the nova subsection; I. Sadeh, simulations and drafting of the capabilities of CTA for transient detection section; O. Sergijenko, simulations and drafting of the detectability of transients of unknown origin section; L. Sidoli, drafting of the SFXT subsection; A. Spolon, simulations and drafting of the tMSPs subsection; L. Zampieri, simulations of the tMSPs subsection.

### v2

@@ 5 Synergies with large astronomical facilities
Simultaneous coordinated observations with telescopes and facilities at different wavelengths are crucial for understanding the processes and mechanisms at work in the sources of our interest (see Cherenkov Telescope Array Consortium et al., 2019 for a detailed review). Coordination, response to real-time alerts and target-of-opportunity (ToOs) are key in time domain astrophysics when observing transient events. In the case of Galactic transients, external alerts can trigger ToO observations by CTAO of new Galactic events or renewed activity of known sources. The trigger criteria is dependent on the type of source observed, with different time urgency and duration varying in terms of the evolution of the specific phenomena. Regarding the observational strategy, a first description on the triggering criteria, expected trigger rate and target priority is already provided in Cherenkov Telescope Array Consortium et al. (2019). Updated strategies for multi-wavelength follow-up and needs are already being collected and discussed in a dedicated operations requirement document999In preparation, where different studies for the planning and optimization of the CTAO follow-up programme are being conducted.
For a prompt reaction and fast reposition, CTAO will need to manage external and internal scientific alerts, for which a transient handler subsystem is under development (Egberts et al., 2022). The LST-1 transient handler is already operational and has allowed the follow-up of different transient alerts (Carosi et al., 2021). Furthermore, a real-time analysis will be performed by CTAO via the Science Alert Generation system with the goal of issuing science alerts of transient phenomena to the community with a latency of about 20 seconds (Bulgarelli et al., 2024), increasing the synergies of CTAO with the astronomical community.
@@
The radio band is of general interest for providing information of the non-thermal processes, and radio facilities are key for locating acceleration sites and shocks and to provide feedback on fast events such as magnetar flares or stellar flares. For that purpose, CTAO will need an external trigger from a radio observatory sensitive to (millisecond) bursts and capable of issuing prompt alerts. Current generation of telescopes such as CHIME, with a large field of view and collecting area, have proven their efficiency for detecting these alerts (CHIME/FRB Collaboration et al., 2020). An example of clear synergies are those with the MeerTRAP project to follow-up fast radio signals (Rajwade et al., 2020; Bezuidenhout et al., 2022)
@@
In the optical and infrared domain prompt reaction would be facilitated by coordinating with external observatories within time zones close to those of CTAO sites. In the case of CTAO North, it should be pursuing synergies mainly with La Palma, Tenerife (and CAHA) telescopes in Spain, while in the CTAO South the most appropriate choice would be the different ESO telescopes and other large facilities in Chile, such as the Rubin Observatory Legacy Survey of Space and Time (LSST), which will be key for time-domain astronomy (see Hambleton et al., 2022). The availability of imaging and spectroscopic instrumentation in both hemispheres does largely fulfill the observational needs required for coordinated campaigns. Specific target of opportunity proposals could be eventually placed by the CTAO community. In parallel, small optical telescopes either on-site or operated by nearby CTAO institutions could also play a key role in this effort if the targets are bright enough as in the case of novae or gamma-ray binaries. This could serve to mitigate the often high-time pressure on large telescopes. Additionally, support optical telescopes would ensure a fast follow-up on transients and would provide sufficient coverage in the case of bright sources.
@@
Space missions will allow to study the X-ray domain with improved capabilities. The Space Variable Objects Monitor (SVOM) [4 - 150 keV range] has been recently launched and although its core science is focused in GRB detection, it has the capabilities for detecting i.e. thermonuclear explosions and outbursts of Galactic origin due to the sensitivity of the ECLAIRs (50mCrab per orbit ) telescope (Wei et al., 2016). Synergies with future instruments such as the Advanced Telescope for High-ENergy Astrophysics (Athena) [0.2–12 keV] (Barcons et al., 2017), scheduled for launch in mid 2030s contemporaneous to the CTAO era, are also expected.
@@
The high-energy band (E100 MeV) is currently explored by the Fermi-LAT satellite (Atwood et al., 2009). Triggers from Fermi-LAT have been important for the detection of Galactic transients such as novae in the VHE domain. Simultaneous MeV-GeV information is also important to disentangle between hadronic and leptonic processes. If Fermi lifetime extends through (at least partially) CTAO lifetime, it will provide the needed coverage and alerts to certain observations/triggers. Just recently, the AGILE satellite, that also operated in this energy range, ceased its operations and re-entered the atmosphere (Tavani et al., 2024). For the moment, there is no obvious successor to Fermi or AGILE in the same energy domain, although some missions could take over. A successor will be crucial to fully exploit the full potential of gamma-ray astronomy. The High Energy cosmic-Radiation Detection facility (HERD) is expected to be operational from 2026 on board of the Chinese Space Station, detecting gamma rays above 0.5 GeV and with potential for transient detection (Cagnoli et al., 2024).
@@
Finally, in a multi-messenger context, CTAO will be able to search for the electromagnetic counterpart of gravitational waves (GWs) and neutrinos in the Milky Way. The Laser Interferometer Space Antenna (LISA) mission will be the first GW observatory on space (scheduled launch on 2030) and will key to study white dwarf mergers (Lamberts et al., 2019; Georgousi et al., 2023). Neutrino alerts from the next generation of detectors such as IceCube-Gen2 will also be crucial in the case of a Galactic core-collapse supernova (Aartsen et al., 2021).
@@ 6 Summary and conclusions
We have studied the capabilities of the forthcoming CTAO to detect transient and variable emission from Galactic sources of different nature. We additionally tested longer exposure times in specific cases of high interest for the observatory. CTAO will be able to discover new transients with not significant degradation in the sensitivity, with a maximum of 15% in the crowded inner regions of the Galaxy when overlapping with strong emitters. Similarly, in order to detect variability from dim systems, our simulations have shown that sources with a photon flux ph cm-2 s-1 will require h of observations to detect this variability. For sources with fluxes above this threshold, only h are required. In the case of strong sources ( ph cm-2 s-1), short exposures h are required, implying that low variations in the flux can be detectable from bright sources. For the case of generic transient sources with the fluxes ph cm-2 s-1 TeV-1 CTAO will not detect any source in about one hour observation time, while for those with fluxes ph cm-2 s-1 TeV-1 in an uncrowded region, CTAO would be able to detect of sources within just 1 hour of observation for both arrays for the different zenith angles and configurations of the geomagnetic field used in this study if all sources are visible. Any visibility constraints will significantly lower this number.
The unique sensitivity at short timescales together with the fast slewing capabilities of the LSTs and the aforementioned capacities of CTAO will allow the detection and discovery of a variety of sources of different nature, according to our simulations:
@@
CTAO will detect VHE from microquasars and from the interaction between their jets and the surrounding environment. Our simulations show that CTAO will likely detect both transient and persistent emission from the massive microquasars Cyg X-1 and Cyg X-3. CTAO will also significantly detect SS 433 including possible flux variability. In the case of LMXBs, CTAO will detect outbursts within few tens of minutes from a nearby source (4 kpc) with relatively small inclination angle (30∘). Longer exposure times are required for LMXBs with larger angles.
@@
We tested the case of tMSPs, concluding that CTAO will need long integration times (50 h) to be able to detect the possible emission of tMSPs when they are in the LMXB state. These systems could be detected during a transition from RSMP to LMXB if an additional VHE component is present, which could provide crucial information on particle interaction.
@@
Flaring emission from the Crab Nebula will be best detected by CTAO (or LST sub-array) at low energies (E200 GeV) in less than 1 h. In the TeV regime, integration times of 10 h will be needed, specially for the detection of dimmer flares.
@@
In the case of novae, CTAO will be able to detect close-by novae. As an example of the only VHE novae known to date, our simulations reveal that CTAO will detect the symbiotic recurrent nova RS Oph with high significance in only 30 min, allowing for a detailed measurement, and therefore detailed modeling, of its SED from energies as low as 20 GeV. Combined with multi-wavelength observations, the temporal and spectral analyses of CTAO observation would improve our understanding of the acceleration processes in novae.
Regarding sources of different nature from the aforementioned, we could expect CTAO to detect emission from magnetars during a giant flare and even likely during intermediate flares associated with an FRB. Other possible transient events are flares from SFXTs, for which a detection would definitely identify SFXTs as VHE emitters. Other variable VHE candidates are runaway stars and young stellar objects. Serendipitous discoveries are also possible while performing, e.g., surveys. Simultaneous multi-wavelength and multi-messenger observations will be crucial to maximize the scientific output of the CTAO observatory.
@@ Acknowledgements
,101010https://www.gammapy.org a community-developed core Python package for TeV gamma-ray astronomy.
@@
The official CTAO observatory IRFs (Maier et al., 2023) used in this manuscript are available in: prod3b-v2
@@
A. Aguasca-Cabot, simulations and drafting of the novae subsection; M. Chernyakova, C. Dignam, simulations and drafting of the SS 433 and microquasars subsections; D. Kantzas, simulations and drafting of the LMXBs subsection and modeling of Cygnus X-1 microquasar; A. López-Oramas, coordinator and editor of the manuscript, general drafting; P. L. Luque-Escamilla, drafting of optical and infrared synergies; S. Markoff, simulations and drafting of the LMXBs subsection and modeling of Cygnus X-1 microquasar; J. Martí, drafting of optical and infrared synergies; D. de Martino, internal review of the manuscript; S. McKeague, simulations and drafting of the SS 433 and microquasars subsections; S. Mereghetti, drafting of the magnetar subsection; E. Mestre, simulations and drafting of the Crab Nebula flares subsection; A. Mitchell, simulations and drafting of the novae subsection; E. de Oña-Wilhelmi, simulations and drafting of the Crab Nebula flares and the novae subsections; G. Piano, simulations and drafting of the Cygnus X-3, Cygnus X-1, V404 Cyg and microquasars subsections; P. Romano, early perspectives with CTAO, SFXT discussion and internal review of the manuscript; A. J. Ruiter, simulations and drafting of the nova subsection; I. Sadeh, simulations and drafting of the capabilities of CTAO for transient detection section; O. Sergijenko, simulations and drafting of the detectability of transients of unknown origin section; L. Sidoli, SFXTs and other sources; A. Spolon, simulations and drafting of the tMSPs subsection; L. Zampieri, simulations of the tMSPs subsection.